# Lipschitz-free spaces and Bossard’s reduction argument

### v3

@@ 2. Notation and preliminaries
Bossard coded the class of separable infinite-dimensional Banach spaces as a $\mathbf{\Delta}^{1}_{1}$ subset $\mathcal{SB}$ of the set of closed subsets of $C(2^{\omega})$ equipped with the Effros-Borel structure. Following a programme initiated in [19], the authors of [10, 9] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$. First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ whose canonical extension to $c_{00}$ is a norm. Endowed with the pointwise topology, $\mathcal{B}$ is a $G_{\delta}$ in $0,\infty)^{V}$ and is thus Polish. Every separable and infinite-dimensional Banach space $(X,\left\|\cdot\right\|)$ has representatives in $\mathcal{B}$. Given a dense linearly independent sequence $(x_{i})\subseteq X$, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$. Then $X\equiv X_{\mu}$, where $X_{\mu}$ is the completion of $(V,\mu)$; conversely, every such completion is a separable and infinite-dimensional Banach space. We will work with $\mathcal{B}$ instead of $\mathcal{SB}$ because, in our opinion, it is easier to use and more compatible with $\mathcal{R}$. While not directly relevant to this paper, another advantage of using $\mathcal{B}$ is that Baire category arguments can be applied, e.g. the set of $\mu\in\mathcal{B}$ such that $X_{\mu}\equiv\mathbb{G}$, where $\mathbb{G}$ is the Gurariĭ space, is a dense $G_{\delta}$ [9]*Theorem 4.1, hence $\mathbb{G}$ can be thought of as the generic separable Banach space. According to [10]*Theorem 8, there is a Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ such that $\Theta(\mu)\equiv X_{\mu}$ for all $\mu\in\mathcal{B}$. It follows easily from this and [6]*Theorem 2.3 (ii) that, given any separable Banach space $X$, the set $\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow}:=\left\{{\mu\in\mathcal{B}}\;:\;\,{X_{\mu}\hookrightarrow X}\right\}$ is $\mathbf{\Sigma}^{1}_{1}$.
@@
Concerning Polish versus Effros-Borel structures, we should mention that it is possible to code the class of complete infinite separable metric spaces using the set of closed infinite subsets of the Urysohn universal separable metric space $\mathbb{U}$, again equipped with the Effros-Borel structure, see e.g. [15, 14]. However we shall work with Vershik’s space $\mathcal{R}$ for the reasons given above. Also, analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [30]*Theorems 1 and 3. Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [13], could be regarded as the generic separable free space. A immediate consequence of this and [13]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual. This observation complements Theorem 4.1.
@@ 3. Three main results
The map $\Xi$ is continuous. Indeed, let $d\in\mathcal{R}$ and $v_{1},\ldots,v_{k}\in V$. There exists $N\in\omega$ such that $v_{\ell}(i)=0$ whenever $\ell=1,\ldots,k$ and $i\geqslant N$. Consider any $d^{\prime}\in\mathcal{R}$ such that $d^{\prime}(i,j)=d(i,j)$ whenever $i,j\leqslant N$. Then $\Xi(d^{\prime})(v_{\ell})=\Xi(d)(v_{\ell})$ for $\ell=1,\ldots,k$. This holds because, given $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$, by McShane’s extension theorem [31]*Theorem 1.33 there exists $g\in B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$, and vice-versa. Therefore $\Xi$ is continuous. ∎
@@
Let $M=(M,\theta)$ be a separable metric space, let $(x_{i})_{i\in\omega}$ be a dense sequence in $M$ whose elements repeat infinitely often, let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.1) and define a second metric $\eta$ on $\omega^{<\omega}$ by
@@
Let $\eta$ be the metric from Lemma 3.7, where $(M,\theta)$ is arbitrary and $\rho$ is as above. Then $(\omega^{<\omega},\eta)$ is discrete, and in addition $(T,\eta)$ is complete whenever $T\in\mathbf{WF}$.
@@
Let $M=(M,\theta)$ be a separable metric space. There exists a continuous map ${\Delta}:{\mathbf{Tr}}\to{\mathcal{R}}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$ whenever $T\in\mathbf{WF}$ and $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$ whenever $T\in\mathbf{IF}$.
@@
Let $(\omega^{<\omega},\eta)$ be the metric space obtained from Lemma 3.7 applied to $M$. Let $W=\omega^{<\omega}\cup(\omega^{<\omega}\cup\{0\})$. We extend $\eta$ to a metric on $W$, again labelled $\eta$, by setting
@@
Let $M$ be a complete separable metric space. By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. By Lemma 3.10, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$. By considering e.g. $\mathbb{U}$ or $c_{0}$ (the latter being Lipschitz universal by Aharoni’s theorem, see e.g. [5]*Theorem 7.11), similarly to above we see that $M$ is also Lipschitz universal. ∎

### v4

@@ 2. Notation and preliminaries
Bossard coded the class of separable infinite-dimensional Banach spaces as a $\mathbf{\Delta}^{1}_{1}$ subset $\mathcal{SB}$ of the set of closed subsets of $C(2^{\omega})$ equipped with the Effros-Borel structure. Following a programme initiated in [18], the authors of [9, 8] (and related papers) observed that it is possible to code the class of separable infinite-dimensional Banach spaces using an arguably canonical Polish space $\mathcal{B}$. First they let $V$ be the countable $\mathbb{Q}$-linear vector space of elements of $c_{00}$ having rational entries, and then they define $\mathcal{B}$ to be the set of all functions $\mu\in[0,\infty)^{V}$ which can be extended (uniquely) to a norm on $c_{00}$. Endowed with the pointwise topology, $\mathcal{B}$ is a $G_{\delta}$ in $0,\infty)^{V}$ and is thus Polish. Every separable and infinite-dimensional Banach space $(X,\left\|\cdot\right\|)$ has representatives in $\mathcal{B}$. Given a linearly independent sequence $(x_{i})\subseteq X$ having dense span, define $\mu(v)=\left\|{\sum_{i\in\omega}v(i)x_{i}}\right\|$, where $v\in V$. Then $X\equiv X_{\mu}$, where $X_{\mu}$ is the completion of $(V,\mu)$; conversely, every such completion is a separable and infinite-dimensional Banach space. We will work with $\mathcal{B}$ instead of $\mathcal{SB}$ because, in our opinion, it is easier to use and more compatible with $\mathcal{R}$. While not directly relevant to this paper, another advantage of using $\mathcal{B}$ is that Baire category arguments can be applied, e.g. the set of $\mu\in\mathcal{B}$ such that $X_{\mu}\equiv\mathbb{G}$, where $\mathbb{G}$ is the Gurariĭ space, is a dense $G_{\delta}$ [8]*Theorem 4.1, hence $\mathbb{G}$ can be thought of as the generic separable Banach space. According to [9]*Theorem 8, there is a Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ such that $\Theta(\mu)\equiv X_{\mu}$ for all $\mu\in\mathcal{B}$. It follows easily from this and [5]*Theorem 2.3 (ii) that, given any separable Banach space $X$, the set $\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow}:=\left\{{\mu\in\mathcal{B}}\;:\;\,{X_{\mu}\hookrightarrow X}\right\}$ is $\mathbf{\Sigma}^{1}_{1}$.
@@
Concerning Polish versus Effros-Borel structures, we should mention that it is possible to code the class of complete infinite separable metric spaces using the set of closed infinite subsets of the Urysohn universal separable metric space $\mathbb{U}$, again equipped with the Effros-Borel structure, see e.g. [14, 13]. However we shall work with Vershik’s space $\mathcal{R}$ for the reasons given above. Also, analogously to $\mathbb{G}$, the set of $d\in\mathcal{R}$ such that $M_{d}\equiv\mathbb{U}$ is a dense $G_{\delta}$ [29]*Theorems 1 and 3. Thus $\mathcal{F}({\mathbb{U}})$, called the Holmes space in [12], could be regarded as the generic separable free space. An immediate consequence of this and [12]*Theorem 2.1 is that the set of $d\in\mathcal{R}$ for which $\mathcal{F}({M_{d}})$ has the metric approximation property is residual. This observation complements Theorem 4.1.
@@ 3. Three main results
It remains to show that $\Xi$ is continuous. Let $d\in\mathcal{R}$, $v_{1},\ldots,v_{k}\in V$ and $\varepsilon\in(0,1)$. To establish the continuity of $\Xi$, we shall find $N\in\omega$ and $\delta>0$ such that
$|\Xi(d^{\prime})(v_{\ell})-\Xi(d)(v_{\ell})|<\varepsilon,\quad 1\leqslant\ell\leqslant k,$
whenever $d^{\prime}\in\mathcal{R}$ satisfies $|d^{\prime}(i,j)-d(i,j)|<\delta$ for all $i,j\leqslant N$.
To begin, we pick $N\in\omega$ such that $v_{\ell}(i)=0$ whenever $1\leqslant\ell\leqslant k$ and $i\geqslant N$. Then set
$r=\min\left\{{d(i,j)}\;:\;\,{i<j\leqslant N}\right\}>0,\quad L=1+\max\left\{{\Xi(d)(v_{\ell})}\;:\;\,{1\leqslant\ell\leqslant k}\right\},$
$\alpha=\varepsilon/(L+1)$ and $\delta=\alpha r$.
Now consider any $d^{\prime}\in\mathcal{R}$ such that $|d^{\prime}(i,j)-d(i,j)|<\delta$ whenever $i,j\leqslant N$. Let $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$. Given that
$(1-\alpha)d(i,j)\leqslant d(i,j)-\delta<d^{\prime}(i,j)<d(i,j)+\delta\leqslant(1+\alpha)d(i,j),\quad i<j\leqslant N,$ (3.1)
the restriction of $f$ to $N+1$ is $(1+\alpha)$-$d$-Lipschitz. By McShane’s extension theorem [30]*Theorem 1.33, there exists $g\in(1+\alpha)B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$. It follows that
$\sum_{i\in\omega}v_{\ell}(i)f(i+1)=\sum_{i<N}v_{\ell}(i)f(i+1)$ $=\sum_{i<N}v_{\ell}(i)g(i+1)$
$=\sum_{i\in\omega}v_{\ell}(i)g(i+1)\leqslant(1+\alpha)\Xi(d)(v_{\ell}),\quad 1\leqslant\ell\leqslant k.$
As these inequalities hold for all such $f$, we obtain
$\Xi(d^{\prime})(v_{\ell})\leqslant(1+\alpha)\Xi(d)(v_{\ell})<\Xi(d)(v_{\ell})+\varepsilon<L,\quad 1\leqslant\ell\leqslant k.$
Now we swap $d$ and $d^{\prime}$ and largely repeat the above. Given $f\in B_{\operatorname{Lip}_{0}(M_{d})}$, by (3.1) and McShane’s theorem, we can find $g\in(1-\alpha)^{-1}B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$. It follows that
$\Xi(d)(v_{\ell})\leqslant\frac{1}{1-\alpha}\Xi(d^{\prime})(v_{\ell})=\bigg(1+\frac{\varepsilon}{L}\bigg)\Xi(d^{\prime})(v_{\ell})<\Xi(d^{\prime})(v_{\ell})+\varepsilon,\quad 1\leqslant\ell\leqslant k.$
This completes the proof. ∎
@@
Let $M=(M,\theta)$ be a complete separable metric space, let $(x_{i})_{i\in\omega}$ be a dense sequence in $M$ whose elements repeat infinitely often, let $\rho$ be a metric on $\omega^{<\omega}$ satisfying (3.2) and define a second metric $\eta$ on $\omega^{<\omega}$ by
@@
Let $\eta$ be the metric from Lemma 3.7, where $(M,\theta)$ is an arbitrary separable metric space and $\rho$ is as above. Then $(\omega^{<\omega},\eta)$ is discrete, and in addition $(T,\eta)$ is complete whenever $T\in\mathbf{WF}$.
@@
Let $M=(M,\theta)$ be a complete separable metric space. There exists a continuous map ${\Delta}:{\mathbf{Tr}}\to{\mathcal{R}}$ such that $\Delta(T)\in\mathcal{R}_{\mathrm{dis}}$ whenever $T\in\mathbf{WF}$ and $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$ whenever $T\in\mathbf{IF}$.
@@
Let $(\omega^{<\omega},\eta)$ be the metric space obtained from Lemma 3.7 applied to $M$. Let $W=\omega^{<\omega}\cup(\omega^{<\omega}\times\{0\})$. We extend $\eta$ to a metric on $W$, again labelled $\eta$, by setting
@@
Let $M$ be a complete separable metric space. By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$. By Lemma 3.10, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$. ∎